There was an equal number of boys and girls in a school auditorium. After 24 boys left the auditorium, there were 3 times as many girls as boys. How many children were there in the beginning?

Answer by math_helper(2461)


72
----------------
Workout:
B = G (1)
3(B-24) = G (2)
Eq (1) allows us to substitute B for G on RHS of (2):
3(B-24) = B
2B = 72
B = 36 ==> G = 36 and B+G=72

Check: 3(B-24) = 3(12) = 36 = G (ok)